# LLM-Advisor: An LLM Benchmark for Cost-efficient Path Planning across Multiple Terrains

### v1

@@ III-A Experimental Setup and Evaluation Metrics
We utilize a GPT-4o to implement our proposed prompt method. We applied the proposed LLM-Advisor to check the planned path generated by A* [13], RRT* [22], and LLM-A* [34]. A* is a search-based method, RRT* is a sampling-based method, and LLM-A* is an LLM-based method. Note that the experiments with LLM-A* [34] also use a GPT-4o. To further evaluate the effectiveness of the LLM-Advisor, we also compare it with LLM-A* [34] and GPT-4o.
Overview of A*. A* is a search-based path planning algorithm. It evaluates the eight neighboring points around the current position and determines the most cost-efficient path by selecting the next step with the lowest combined cost of traversal and estimated remaining distance(Fig. 6). Therefore, A* generates reasonable paths, as illustrated in Fig. 7. However, the generated path often includes unnecessary detours. Although A* avoids obstacles, the overall path may not be the most cost-efficient.
TABLE III: Evaluation of planned paths across different methods on the MultiTerraPath dataset using the proposed LLM-Advisor
Method Terrain description Path description $N_{\text{Images}}$ $N_{\text{Path}}$ $N_{\text{Suggested}}$ $N_{\text{Improved}}$ $RP$ (%) $\uparrow$
@@
LLM-Advisor + A* [13] - 1,000 1,000 612 432 70.59
LLM-Advisor + RRT* [22] - 1,000 1,000 737 512 69.47
LLM-Advisor + LLM-A* [34] - 1,000 1,000 770 606 78.70
TABLE IV: Evaluation of planned paths across different methods on the RUGD_v2 dataset using the proposed LLM-Advisor
Method Scenes $N_{\text{Images}}$ $N_{\text{Path}}$ $N_{\text{Suggested}}$ $N_{\text{Improved}}$ $RP$ (%) $\uparrow$
LLM-Advisor + A* Creek 1000 1000 876 580 66.21
Park 1000 1000 860 720 83.72
Trail 1000 1000 620 540 87.10
Village 1000 1000 780 560 71.79
TABLE V: Comparison of our LLM-Advisor with other LLM-based methods
Method Terrain description $N_{\text{Images}}$ $N_{\text{Path}}$ $N_{\text{Successful}}$ $N_{\text{Improved}}$ $N_{\text{Equal}}$ $N_{\text{Deteriorated}}$ $IR$ (%) $\uparrow$ FPS $\uparrow$
A* [13] - 1,000 1,000 1,000 - - - - 9.39
GPT-4o 1,000 1,000 1,000 208 0 792 20.80 1.21
LLM-A* [34] 1,000 1,000 938 152 538 248 69.00 3.10
LLM-Advisor + A* 1,000 1,000 1,000 432 393 175 82.50 5.86
Figure 9: Visualization of the planned paths by GPT-4o, A*, LLM-A*, and LLM-Advisor + A*. Black bold numbers indicate the favorable suggestions, while red bold numbers highlight less favorable suggestions. All other numbers represent cases where no suggestions were made. Lower costs indicate better performance
Figure 10: Visualization of experimental results on RUGD_v2 dataset
TABLE VI: Ablation studies
Method Terrain description Path description RAG Descriptive output $N_{\text{Images}}$ $N_{\text{Path}}$ $N_{\text{Suggested}}$ $N_{\text{Improved}}$ $RP$ (%) $\uparrow$
Detailed Brief
LLM-Advisor + A* - - - 1,000 1,000 602 130 21.59
- - - 1,000 1,000 607 352 57.99
- - 1,000 1,000 610 413 67.70
- - 1,000 1,000 609 399 65.52
- 1,000 1,000 612 432 70.59
Path description (Brief): From (88, 172) to (305, 370) the terrain cost is 1. From (306, 371) to (306, 371) the terrain cost is 9223372036854775807. From (307, 372) to (337, 413) the terrain cost is 1. From (337, 414) to (337, 443) the terrain cost is 5. The total cost of the path is -9223372036854775399.
Figure 11: Path description (Brief)
Figure 12: Visualization of experimental results after integrating the proposed hallucination-mitigation strategies
Overview of RRT*. RRT* is a sampling-based path planning algorithm that extends RRT [21]. It randomly samples points and connects each new point to the nearest neighbor in the tree. Unlike RRT, RRT* evaluates the cost from all nearby nodes and selects the lowest-cost connection. After connecting, it rewires the tree whenever a more efficient path is found. Over time, this iterative rewiring process improves the overall path quality. Although time-consuming, RRT* generates reasonable paths, which is why we use our LLM-Advisor to evaluate its results, aiming to assess its effectiveness as a validation tool for precomputed reasonable paths.
Overview of LLM-A*. LLM-A* leverages LLMs to provide a candidate list for the A* algorithm. The main idea of LLM-A* is to improve search efficiency in obstacle-avoidance path planning. By restricting the search to the candidate list provided by LLM, LLM-A* can significantly decrease the search time. The prompt used for experiments using LLM-A* [34] is given in Fig. 8.
Since LLM-Advisor functions as an advisor, we assess its performance using two metrics: Relative Precision ($RP$) and Improvement Ratio ($IR$). The definition of $RP$ is provided in Eq. 1, and the definition of $IR$ is given in Eq. 2. $RP$ represents the proportion of improved paths within the suggestions made by LLM-Advisor, while $IR$ indicates the percentage of paths that are equal to or better than the original ones after incorporating LLM-Advisor. $IR$ can reveal how many paths are valid and effective.
@@ III-B Main Results
Main Comparisons. Table III presents the main comparison results for the MultiTerraPath dataset. Most suggestions made by the proposed LLM-Advisor are more cost-efficient. It achieves an $RP$ of 70.59% when making suggestions for the A*-planned path, 69.47% for the RRT*-planned path, and 78.70% for the LLM-A*-planned path. This demonstrates that, with detailed path descriptions, LLM-Advisor can effectively serve as a reliable validation tool for planned paths generated by other methods. Moreover, the proposed LLM-Advisor is versatile. It can improve the previous approaches regardless of whether they are search-based, sampling-based, or LLM-based. Table IV shows experimental results for the the RUGD_v2 dataset.
@@
Table V presents a comparison of LLM-Advisor+A* with A*, LLM-A* [34] and GPT-4o. The best performance for each evaluation metric is bolded, while the second-best is underlined. A* successfully generated valid paths for all cases. Since the proposed LLM-Advisor functions as an advisor, evaluating the paths generated by A*, it also successfully validated all paths. GPT-4o successfully planned all paths, although most were suboptimal or invalid. However, it was still able to generate a path. In LLM-A [34], GPT-4o suggests a target list, providing a sequence of intermediate target states for A* to consider during path planning. If the generated target list contains points within the obstacle region, the neighbor search process cannot proceed. As shown in Table V, there were 62 paths that were not successfully planned using LLM-A*. When considering the Improvement Ratio (IR), LLM-Advisor+A* achieved a higher IR, indicating better performance. The GPT-4o has the lowest IR, indicating that GPT-4o does not have ability to do zero-shot path planning. Moreover, since our LLM-Advisor acts as a postprocessor, it is possible to involve humans in the verification process, allowing us to decide whether to accept the paths suggested by the LLM-Advisor. Incorporating human judgment in this process can further improve the quality of the planned paths.
To evaluate the computational efficiency of different path planning methods, we report frames per second (FPS) for each approach in Table V. FPS quantifies the number of paths processed per second, where a higher value indicates greater efficiency. All methods are executed entirely on the CPU. Among them, LLM-Advisor + A* selectively makes suggestions only when an improvement is needed, minimizing unnecessary LLM computations and achieving a higher FPS compared to other LLM-based methods.
Visualizations. Fig. 9 presents some path planning results. As shown in the figure, our LLM-Advisor successfully suggests more cost-efficient paths. Fig. 10 provides visualization results for the RUGD_v2 dataset.
@@
Ablation Studies. Table VI shows some ablation study results. We evaluate LLM-Advisor with both brief and detailed path descriptions to demonstrate the effectiveness of the proposed prompt. One example of brief path description is given in Fig. 11. As shown, the level of detail in path descriptions significantly impacts final performance. By using detailed path descriptions, the LLM-Advisor can more accurately evaluate the A*-planned path and make a higher-quality suggestion. We also found that it has only a limited influence on the number of suggestions. This indicates that the ability of LLM-Advisor to suggest improvements to planned paths is not affected much by the path descriptions. Moreover, using the proposed hallucination-mitigation strategies, the LLM-Advisor achieves higher RP. Fig. 12 shows the visualization results.
@@ IV Future Work
The proposed LLM-Advisor operates as a zero-shot model, relying heavily on the performance of general-purpose pretrained LLMs. Without task-specific training, it leverages the inherent knowledge and reasoning capabilities of these models to provide guidance and insights. However, its performance may be limited when dealing with domain-specific challenges. As a next step, we plan to fine-tune the model using task-relevant data to enhance its accuracy, adaptability, and effectiveness in path planning and related tasks.
@@ V Conclusions
This paper introduced a novel prompt-based approach, LLM-Advisor, which achieves high relative precision and improvement ratio in evaluating paths planned by existing methods. The LLM-Advisor is versatile and can be used to verify paths generated by any planing method. It provides suggestions when existing methods fail to identify the most cost-efficient path. Furthermore, we propose two strategies to mitigate hallucinations. Our experiments on created MultiTerraPath and RUGD_v2 datasets demonstrate the effectiveness of LLM-Advisor. We also experimentally found that GPT-4o struggles with effective path planning based solely on terrain descriptions. We believe that our work offers valuable insights for the path planning community, particularly for researchers investigating the integration of LLMs into path planning.
Acknowledgments
This paper is partially financially supported by JSPS KAKENHI Grant Number 24K20787, Fund to Support Female Researchers with Their Research Skills Training of The University of Tokyo, and Institute for AI and Beyond of The University of Tokyo.

### v2

@@ III-A Experimental Setup and Evaluation Metrics
We applied the proposed LLM-Advisor to check the planned path generated by A* [4], RRT* [7], and LLM-A* [16]. A* is a search-based method, RRT* is a sampling-based method, and LLM-A* is an LLM-based method. The prompt used for LLM-A* [16] is given in Fig. 7. To further evaluate the effectiveness of LLM-Advisor, we also compare it with LLM-A* [16], GPT-4o, GPT-4-turbo, Gemini-2.5-flash, and Claude-Opus-4. We adopt the same cost calculation method as A* [4] for all compared approaches.
TABLE III: Evaluation of planned paths across different methods on the MultiTerraPath dataset using the proposed LLM-Advisor.
Method Terrain description Path description Easy subset Hard subset
Detailed Brief RP (%) $\uparrow$ SR (%) $\uparrow$ SCP (%) $\uparrow$ RP (%) $\uparrow$ SR (%) $\uparrow$ SCP (%) $\uparrow$
LLM-Advisor (GPT-4o) + A* [4] - 72.37 99.10 94.10 74.56 99.30 94.32
LLM-Advisor (GPT-4o) + RRT* [7] - 69.47 96.00 88.25 73.21 96.70 91.58
LLM-Advisor (GPT-4o) + LLM-A* [16] - 78.70 97.00 84.47 81.39 97.70 89.91
TABLE IV: Evaluations on the RUGD_v2 dataset using the proposed LLM-Advisor. $N_{\text{Images}}=1,000$ and $N_{\text{Path}}=1,000$.
Method Scenes $N_{\text{Suggested}}$ $N_{\text{Improved}}$ RP (%) $\uparrow$ SR(%) $\uparrow$ SCP(%) $\uparrow$
LLM-Advisor (GPT-4o) + A* Creek 876 580 66.21 94.70 84.39
Park 860 720 83.72 99.10 89.11
Trail 620 540 87.10 100.00 95.23
Village 780 560 71.79 96.20 86.82
TABLE V: Comparisons with other LLM-based methods, evaluated on easy subset of MultiTerraPath. $N_{\text{Images}}=1,000$ and $N_{\text{Path}}=1,000$.
Method Terrain description $N_{\text{Improved}}$ $N_{\text{Equal}}$ $N_{\text{Deteriorated}}$ IR (%) $\uparrow$ SR(%) $\uparrow$ SCP(%) $\uparrow$ FPS $\uparrow$
A* [4] (Baseline) - - - - - 100.00 86.02 9.39
GPT-4o 208 0 792 20.80 43.10 31.23 0.08
GPT-4-turbo 233 0 767 23.30 46.20 34.00 0.09
Gemini-2.5-flash 48 0 952 4.80 27.10 20.26 0.03
Claude-Opus-4 207 0 793 20.70 41.00 29.43 0.08
LLM-A* [16] 152 538 248 69.00 93.80 87.29 0.81
LLM-Advisor (GPT-4o) + A* 626 135 239 76.10 99.10 94.10 1.09
LLM-Advisor (GPT-4-turbo) + A* 268 495 237 76.30 99.20 95.23 1.26
TABLE VI: Ablation studies when using easy subset of MultiTerraPath. $N_{\text{Images}}=1,000$ and $N_{\text{Path}}=1,000$. As more structured context and grounding are introduced, LLM-Advisor produces increasingly precise and reliable suggestions, while hallucination rates consistently decrease.
Method Terrain description Path description RAG Descriptive output $N_{\text{Suggested}}$ $N_{\text{Improved}}$ RP (%) $\uparrow$ IR (%) $\uparrow$ HR (%) $\downarrow$
@@
LLM-Advisor (GPT-4o)+ A* - - - 602 130 21.59 52.80 3.30
- - - 607 352 57.99 74.50 2.00
- - 610 413 67.70 80.30 1.10
- - 609 399 65.52 79.00 1.40
- 865 626 72.37 76.10 0.70
Figure 8: Results on MultiTerraPath dataset with LLM-Advisor (GPT-4o) + A*.
Figure 9: Results on RUGD_v2 dataset with LLM-Advisor (GPT-4o) + A*.
Figure 10: Results in challenging environments: (a) detours around trees in a blurry image; (b) robustness to lighting and overexposure.
Path description (Brief): From (88, 172) to (305, 370) the terrain cost is 1. From (306, 371) to (306, 371) the terrain cost is 9223372036854775807. From (307, 372) to (337, 413) the terrain cost is 1. From (337, 414) to (337, 443) the terrain cost is 5. The total cost of the path is 9223372036854775399.
Figure 11: Path description (Brief).
Figure 12: Results on the hard subset of MultiTerraPath. We visualize the original A* paths and the improved paths refined by LLM-Advisor (GPT-4o). LLM-Advisor shows stronger effectiveness in hard scenarios with complex terrains and obstacles.
Figure 13: Results after applying hallucination-mitigation strategies.
We assess the performance of LLM-Advisor using five metrics: Relative Precision (RP), Improvement Ratio (IR), Success Rate (SR), newly proposed cost-aware success weighted by path length (SCP) and hallucination rate (HR).
1) RP represents the proportion of improved paths among all suggestions made by LLM-Advisor. The definition of RP is provided in Eq. 1.
@@
2) IR indicates the percentage of paths that are equal to or better than the original A*-planned paths after incorporating LLM-Advisor. Therefore, IR reflects how many suggested paths are valid and effective. The definition of IR is given in Eq. 2.
@@
3) SR measures the proportion of planned paths that successfully reach the target without collision. It is defined as
$SR=\frac{N_{\text{Successful}}}{N_{\text{Path}}},$ (3)
where $N_{\text{Successful}}$ denotes the number of collision-free paths that reach the goal.
4) SCP evaluates navigation performance based on traversal cost rather than geometric path length. The commonly used Success weighted by Path Length (SPL) metric assumes uniform traversal cost and evaluates path optimality purely based on path length, which becomes inadequate in environments with heterogeneous terrain costs. In such cases, a shorter path may incur higher traversal cost due to unfavorable terrain. To address this limitation, we propose SCP, defined as
$\mathrm{SCP}=\frac{1}{N_{\text{Path}}}\sum_{i=1}^{N_{\text{Path}}}S_{i}\cdot\frac{C_{i}^{\ast}}{\max\left(C_{i},\,C_{i}^{\ast}\right)},$ (4)
where $S_{i}\in\{0,1\}$ is a per-path success indicator, $C_{i}$ denotes the total traversal cost of the path generated by the model, and $C_{i}^{\ast}$ represents the minimum-cost path obtained by a cost-aware shortest-path planner. The traversal cost $C_{i}$ is computed as
$C_{i}=\sum_{k=1}^{K_{i}}d_{k}\cdot w(t_{k}),$ (5)
where $d_{k}$ is the length of the $k$-th path segment, $t_{k}$ denotes the terrain type traversed by the segment, and $w(\cdot)$ is the corresponding terrain-dependent cost weight. By explicitly modeling terrain-dependent traversal cost, SCP generalizes SPL to cost-aware navigation settings and provides a more faithful measure of navigation efficiency in heterogeneous environments.
5) HR measures the proportion of planned paths that are physically invalid or hallucinated. It is defined as
$HR=\frac{N_{\text{Hallucinated}}}{N_{\text{Path}}},$ (6)
where $N_{\text{Hallucinated}}$ denotes the number of paths that violate physical or environmental constraints, such as trajectories that leave the valid navigation space (e.g., paths going into the air).
@@ III-B Main Results
Main Comparisons. Table III presents the main comparison results for the MultiTerraPath dataset. Most suggestions made by the proposed LLM-Advisor (GPT-4o) are more cost-efficient. When evaluated using RP, SR, and SCP, LLM-Advisor (GPT-4o) demonstrates superior effectiveness in harder scenarios. This demonstrates that, with detailed path descriptions, LLM-Advisor can effectively serve as a reliable validation tool for planned paths generated by other methods. Moreover, LLM-Advisor is versatile. It can improve the previous approaches regardless of whether they are search-based, sampling-based, or LLM-based. Table IV shows experimental results for the RUGD_v2 dataset.
@@
Table VI presents a comparison of LLM-Advisor+A* with A*, GPT-4o, GPT-4-turbo, Gemini-2.5-flash, Claude-Opus-4, and LLM-A* [16]. The best performance for each evaluation metric is bolded. In LLM-A [16], GPT-4o suggests a target list, providing a sequence of intermediate target states for A* to consider during path planning. If the generated target list contains points within the obstacle region, the neighbor search process cannot proceed. As shown in Table VI, there were 62 paths that were not successfully planned using LLM-A*, achieving a SR of 93.80%. LLM-Advisor (GPT-4o) + A* yields 626 improved paths and achieves the overall Improvement Ratio (IR) of 76.10% and a SR of 99.10%. LLM-Advisor (GPT-4-turbo) + A* achieves the highest IR. The GPT-4o, GPT-4-turbo, Gemini-2.5-flash, and Claude-Opus-4 have relatively lower IR, indicating that existing LLMs does not have ability to do zero-shot path planning. Moreover, since our LLM-Advisor acts as a postprocessor, it is possible to involve humans in the verification process, allowing us to decide whether to accept the paths suggested by LLM-Advisor. Incorporating human judgment in this process can further improve the quality of the planned paths.
To evaluate computational efficiency, we report frames per second (FPS) for each method in Table VI. All methods run entirely on the CPU. LLM-Advisor + A* achieves a higher FPS than other LLM-based approaches.
In addition to runtime efficiency, we also report token efficiency for LLM-Advisor. Token usage is measured via the OpenAI API statistics (prompt and completion tokens). In our implementation, each inference call corresponds to one start–goal decision and consumes approximately 3.2k tokens, with a negligible completion length when the A* path is already cost-effective.
Visualizations. As shown in Figs. 8 and 10, our LLM-Advisor successfully suggests more cost-efficient paths. Figs. 10 and 13 present experimental results in more challenging environments. Our model demonstrates robustness to visual degradations such as blurry images and overexposure (Fig. 10), and is able to plan significantly more cost-effective paths in harder scenarios (Fig. 13).
@@
Ablation Studies. We experimentally verified the difference between detailed and brief path representations (Fig. 11). As shown in Table VI, the granularity in path descriptions significantly impacts final performance. By using detailed path descriptions, the LLM-Advisor can more accurately evaluate the A*-planned path and make a higher-quality suggestion. We also found that the level of detail has only a limited influence on the number of suggestions. This indicates that the ability of LLM-Advisor to suggest improvements to planned paths is not affected much by the path descriptions. Moreover, using the proposed hallucination-mitigation strategies, the LLM-Advisor achieves higher RP and lower HR. Fig. 13 shows the corresponding visualization results after applying the hallucination-mitigation strategies.
@@ IV Future Work
The proposed LLM-Advisor functions as a zero-shot model, relying heavily on the performance of general-purpose pretrained LLMs. Therefore, its performance may be suboptimal when dealing with domain-specific challenges. As a next step, we plan to fine-tune the model using task-specific data to enhance its accuracy, adaptability, and effectiveness in path planning and related applications. We also aim to extend its capabilities to better handle ambiguous inputs, diverse environments, and noisy terrain descriptions.
@@ V Conclusions
This paper introduces LLM-Advisor, a novel planner-agnostic mechanism for cost refinement that operates without modifying the underlying planner. We further propose two effective strategies to mitigate LLM hallucinations, and introduce the MultiTerraPath and RUGD_v2 datasets for systematic evaluation. Extensive experimental validation demonstrates that: (1) LLM-Advisor is versatile and can be applied to paths generated by a wide range of planning methods, achieving high relative precision and improvement ratio, and providing effective suggestions when existing planners fail to identify the most cost-efficient paths; (2) the proposed hallucination-mitigation strategies significantly reduce hallucination rates; (3) unlike end-to-end planning approaches, LLM-Advisor limits the impact of hallucinations by employing LLMs in a non-decisive advisory role; and (4) state-of-the-art LLMs, including GPT-4o, GPT-4-turbo, Gemini-2.5-Flash, and Claude-Opus-4, struggle to perform effective path planning based solely on terrain descriptions. We believe that this work provides valuable insights for the path planning community, particularly for researchers exploring the integration of LLMs into classical planning frameworks.